OpenStudy (anonymous):

6 men and 8 women can complete a work in 10 days. 26 men and 48 women can finish the same work in 2 days. 15 men and 20 women can do the same work in - days.

OpenStudy (anonymous):

6m+8w=10-->1 26m+48w=2-->2 *eqn 1 into 6 so 36m+48w=60 eqn2 26m+48w=2 m=5.8 ,w=3.1

mathslover (mathslover):

Not exactly the way to solve this problem. The approach will be different.

mathslover (mathslover):

Let us take some common variables to represent the work done by man and woman in 1 day : Work done by 1 man in 1 day : x Work done by 1 woman in 1 day : y

mathslover (mathslover):

So, we can say, that Work done by 6 men in 1 day : 6x Work done by 8 women in 1 day : 8y It is given that 6 men and 8 women complete the work in 10 days . So, 6 men and 8 women will do 1/10 of the total work in 1 day, right?

OpenStudy (anonymous):

yes that i used before then

mathslover (mathslover):

Not exactly. The process you did had a mistake in RHS it should be 1/10 in the first equation in RHS : \(\boxed{\mathsf{ 6x + 8y = \dfrac{1}{10} \quad ------ (1) \\ 26x + 48y = \dfrac{1}{2} \quad ----- (2) \\ }}\)

mathslover (mathslover):

Solve these two equations for x and y and put them into 15 x + 20 y = ? The answer you'll get will be 15 men's and 20 women's 1 day work. Just reciprocal the fraction obtained and you'll get your answer. Good Luck

OpenStudy (anonymous):

x=1/10 ,y=59/80

OpenStudy (anonymous):

not yet

mathslover (mathslover):

Not correct, which method did you use to solve these two equations?

OpenStudy (anonymous):

36x+48y=6/10 26x+48y=1/2

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

6+8y=1/10

mathslover (mathslover):

Okay fine : \(\boxed{\mathsf{\color{blue}{36 x + 48 y = \dfrac{6}{10} \\ 26x + 48 y = \dfrac{1}{2} \\ - \quad - \quad \quad - \\ 10 x + 0 = \dfrac{6}{10 } - \dfrac{1}{2} \\ 10x = \dfrac{6 - 5}{10} \\ 10x = \dfrac{1}{10} \\ }} \color{red}{ \textbf{Divide 10 both sides}} \\ \color{blue}{ \mathsf{\dfrac{10x}{10} = \dfrac{1}{10 \times 10 } \\ \dfrac{\cancel{10} x}{\cancel{10}} = \dfrac{1}{100} \\ x = \dfrac{1}{100} }}}\)